OpenStudy (anonymous):

Which equation could be used to solve the problem? The Jensens drove 400 miles in 8 hours. How many miles per hour did they drive (m)? A. m = 400 miles + 8 hours B. m = 8 hours ÷ 400 miles C. m = 400 miles • 8 hours D. m = 400 miles ÷ 8 hours

OpenStudy (anonymous):

2.Which equation could be used to solve the problem? Oscar averaged 4 yards per carry in the football game. If he carried the ball 16 times, how many yards did he gain (y)? A. y = 4 • 16 B. y = 16 ÷ 4 C. y = 4 ÷ 16 D. y = 4 + 16 The manager of the neighborhood baseball team bought 150 baseballs in boxes of 6. How many boxes did he buy?

OpenStudy (anonymous):

You want to multiply your values in the first one

OpenStudy (anonymous):

@timanti

OpenStudy (anonymous):

basically asking how many m=400x8

OpenStudy (anonymous):

Same as the second one it is needing to be multiplied because of how many times he carried the ball times the amount of yards he went

OpenStudy (anonymous):

the third one is division. 150 balls divided by boxes= the amount of boxes he bought